# Quantum Radiometric Calibration

### v2

@@ Abstract
An entirely different calibration method was presented in [13]. It solely utilises squeezed states of light, the Heisenberg uncertainty principle, and relative photon loss measurements. The detection efficiencies of the presumably best photo diodes on the market for 1064 nm and 1550 nm were calibrated to $(99.5\pm 0.5)$% [13] and $(98.5\pm 0.7)$% [14], respectively, using 15-dB and 13.5-dB squeezed vacuum states of light. (Both calibrations used an angle of incidence of 20, as specified by the manufacturer. In addition, the reflected light, which accounted for approximately 0.5%, was reflected back to the photodiode to maximise detection efficiency.) The method should be suitable for calibrating photodiodes as primary standards, however, the theoretical description of the quality of the calibration signal is still missing. Furthermore, there is currently no precise measurement method for the photon escape efficiency of the squeezing resonator, which represents a potential source of significant systematic error.
In this work, we present the theoretical description of the calibration signal of quantum radiometric calibration (QRC) as well as the required concept for the measurement of the escape efficiency. We apply the concept in-situ and determine the escape efficiency of our squeezing resonator to (98.583 ± 0.015)%. We use “only” 10-dB squeezed vacuum states, but were nonetheless able to precisely determine the efficiency of the presumably most sensitive photodiodes on the market at 1550 nm (from the same manufacturer as calibrated in [14]) to $\eta_{\rm DE}\!=\!(97.20\pm 0.37)$% and $\eta_{\rm QE}\!=\!(96.9\pm 0.4)$% under the specified angle of 20. Implementing retro-reflection increase the values by 0.46%. The value is unexpectedly low and not sufficient for optical quantum computing.
@@
where the voltages $U$ are transimpedance amplified photo currents measured on continuous-wave light. $U_{\mathrm{tot}}$ is the total measurement voltage and $U_{\mathrm{dark}}$ is the contribution without laser light. $U_{\mathrm{perf}}$ represents the perfect case, i.e., the unknown voltage that the measurement system would provide if every photon of the light beam (that is focused on the photodiode surface) would be converted into exactly one photo electron, combined with zero dark current. In the limit of negligible dark noise, $\eta_{\mathrm{QE}}$ approaches $\eta_{\mathrm{DE}}$. The challenge is to determine accurately $\eta_{\mathrm{DE}}$ and thus $\eta_{\mathrm{QE}}$, i.e., to calibrate a photodetector.
@@
According to Eq. 4, $\eta_{\mathrm{DE}}$ is found when $\eta$ and the other three efficiencies are determined. Here, $\eta_{\mathrm{esc}}$ is the photon escape efficiency from the squeezing resonator, $\eta_{\mathrm{prop}}$ the propagation efficiency from the same resonator to the measurement, and $\eta_{\mathrm{mm}}$ the modematching efficiency with respect to the local oscillator beam of the BHD. $\eta_{\mathrm{mm}}$ corresponds to the square of the interference visibility at the BHD beam splitter. $\eta_{\mathrm{prop}}$ corresponds to the transmissivity of the optics along the path from the squeezing resonator to the photodiodes. It is measured with a single photo diode on the auxiliary beam at both ends of the optical path. The squeezing resonator’s photon escape efficiency $\eta_{\mathrm{esc}}$ could previously not be measured but just estimated and turned out to be the by far largest error contribution of the method [13]. The concept for measuring $\eta_{\mathrm{esc}}$ is a result of this work and will be discussed later. The basic principle for measuring $\eta$ in Eq. (4) will be further discussed here first.
The precision increases with an increasing absolute value of the derivative of the measured uncertainty product to $\eta$. The measured uncertainty product is according to Eq. (2) {align} Δ^2 ^X ⋅Δ^2 ^Y = [ η Δ^2 ^X’ + (1-η) ] ⋅ [ η Δ^2 ^Y’ + (1-η) ]
@@
= 1 + 4(η-η^2) ⟨^n ⟩  , where $(\Delta\!^{2}\!\hat{X}^{\prime}+\Delta\!^{2}\hat{Y}^{\prime})/4-0.5=\langle\hat{n}\rangle$ is the expectation value for the number of quantum correlated photons in the ‘Heisenberg reference’ [20]. For the pure 13.2-dB squeezed state in Fig. 1(b) $\langle\hat{n}\rangle\approx 4.7$.
@@
Here, we introduce the ‘precision scaling’ of QRC $|S_{P}|$, which includes the derivative of Eq. (1) to $\eta$ and the fact that the measurement precisions of $\Delta\!^{2}\!\hat{X}$ and $\Delta\!^{2}\hat{Y}$ are both proportional to the total efficiency $\eta$. {align} —S_P— = η^2  — d (1+ 4(η- η2) ⟨^n⟩)d η — = η^2  —4 - 8 η— ⟨^n ⟩  . Thus, the QRC precision increases with $\langle\hat{n}\rangle$ (whereby a limit is set by the fact that the decoherence due to phase fluctuations must remain negligible, see ‘end matter’) and with $\eta_{\mathrm{DE}}$, i.e., with the quality of the photodiode to be calibrated.
The QRC approach offers a very high degree of accuracy, as potential systematic errors in independent measurements can be quantified and minimised to negligible values. All additional measurements are exclusively photoelectric in nature and take place in situ, i.e., in the same laser-optical QRC setup. Another advantage for accuracy is that QRC does not calibrate the direct current of the photodetectors, but rather the spectral power at the sideband frequency of the user application. QRC can provide a comprehensive calibration curve against sideband frequency.
We calibrated at a sideband frequency of 5 MHz a pair of 1550-nm HQE photodiodes from Laser Components (without protection window) optimized for s-polarized continuous-wave laser light at 1550 nm under 20 angle of incidence. The semiconductor surface was free of dust particles. The experiment consisted of a resonator to generate squeezed vacuum states of light at 1550 nm and a BHD for their measurement using two of the photodiodes to be calibrated. The efficiencies $\eta_{\mathrm{esc}}$, $\eta_{\mathrm{prop}}$, and $\eta_{\mathrm{mm}}$ were maximized by selecting high-quality optics and optimizing the modematching aiming for a measured squeeze factor as high as possible. (The three efficiencies were quantified afterwards by separate measurements.) We used the most efficient approach to create squeezed light, which is resonator-enhanced degenerate parametric down-conversion (PDC) in a second-order non-linear crystal [22, 23, 24, 13, 31]. A schematic is shown in Fig. 2.
@@
The crystals in the SHG and squeezing resonators were periodically poled potassium titanyl phosphate (PPKTP). The crystal geometry was plane-convex and formed a half-monolithic resonator together with a coupling mirror. The crystal was coated to be highly reflective on the convex side and anti-reflective on the other for 1550 nm and 775 nm. The coupling mirror had a design reflectivity of $\approx\!83$ % for 1550 nm and $\approx\!97.5$ % for 775 nm light. The temperature profile of the crystal [32, 33, 34] was stabilized to a setpoint that provided double-resonance as well as a good phase matching. The length of the squeezing resonator was stabilized using the Pound-Drever-Hall (PDH) technique [35]. The squeezed vacuum states at 1550 nm without any carrier light ‘escaped’ the squeezing resonator through the coupling mirror, were separated from the 775 nm light via a dichroic mirror, and superimposed with the local oscillator on the BHD’s beam splitter. We directly measured a 10-dB squeezed photon shot noise at a 775-nm pump power of slightly less than 11 mW. The anti-squeezing was close to 13 dB. Increasing the pump power increased the squeeze factor just marginally, while the anti-squeezing increased significantly. We used this measurement to estimate the root mean square (rms) phase noise [36] to $\Delta\theta\!<\!(2\!\pm\!1)$ mrad, which was negligible for the QRC result reported further down.
The escape efficiency $\eta_{\rm esc}$ of the squeezing resonator could previously only be estimated using an absorption value of the crystal material from the literature [37] and rather uncertain crystal and coupling mirror coating specifications from the manufacturer [13]. Furthermore, potential surface inhomogeneities and small dust particles in the squeezing resonator, which make the value of $\eta_{\rm esc}$ dependent on its alignment, could not be taken into account. In [13], the inaccuracy in the determination of $\eta_{\rm esc}$ clearly dominated the calibration error. The escape efficiency of the quantum correlated photons from the squeezing resonator is given by
$\eta_{\rm esc}=\frac{1-r^{2}}{1-r^{2}+\ell_{\mathrm{rt}}^{2}}\,$, (5)
where $r^{2}$ is the reflectivity of the coupling mirror and $\ell_{\mathrm{rt}}^{2}$ is the cavity mode’s relative photon loss per round trip.
Fig. 3 shows our setup for determining $\eta_{\rm esc}$ (left), a measurement graph of the reflected light power (top right), and the simultaneously measured transmitted power (bottom right). The latter provided the modematching value of $(98.58\pm 0.23)$% (main peak divided by sum of all peaks), which was required to scale up the loss dip in the top right graph. The fit of Eq. (6) to the up-scaled top right graph required a simultaneous fit of the scanning speed, which we approximated by $\varphi(t)=\varphi_{0}+\varphi_{1}t+\varphi_{2}t^{2}$. Our in-situ measurement yielded $r^{2}=(82.79\pm 0.35)$% and $\ell^{2}_{\mathrm{rt}}=(0.247\pm 0.007)$% and thus $\eta_{\rm esc}=(98.583\pm 0.015)$% according to Eq. (5).
The propagation efficiency $\eta_{\mathrm{prop}}$ and the mode matching efficiency $\eta_{\mathrm{mm}}$ at the BHD beam splitter were measured in the established manner, which is described in the ‘end matter’ to this letter. The measured values yielded the $\eta_{\mathrm{prop}}=(99.49\,\pm\,0.25)$% and $\eta_{\mathrm{mm}}=v^{2}=(99.11\,\pm\,0.17)\%$.
@@
The key measurement of QRC is that of the total quantum efficiency $\eta$. For this purpose, the variances of the squeezed and anti-squeezed quadratures in Eq. (1) were measured with the BHD. We repeatedly swept the BHD read-out phase $\theta$ to continuously rotate between measuring the maximally squeezed and the maximally anti-squeezed quadratures, see Fig. 4. This approach offers a few advantages over the alternative of locking the read-out phase to each individual quadrature as used before [13, 14]. First of all, it reduces the complexity of the setup by omitting the stabilization of one degree of freedom. Additionally, slow drifts in the local oscillator and pump power can be neglected. Scanning the read-out phase also guarantees that there is no systematic error in the locking point and ensures that each quadrature is optimally measured at some point. The traces in Fig. 4 already include dark noise subtraction and normalisation to the ground state variance. The variance of the ground state was measured with the signal input of the BHD blocked. The variance of the dark noise was measured with all light fields blocked.
@@
The QRC method presented in detail and used here has several unique features compared to established methods. It uses only Heisenberg’s uncertainty principle and the photoelectric effect. No direct current (DC) signals are used for calibration, but rather noise power at relevant measurement frequencies. QRC is suitable for determining quantum efficiencies that depend on the measurement frequency. We have shown that absolute calibration with an uncertainty of only 0.37% is possible with squeezed vacuum states in the achievable 10 dB range. Squeeze lasers of this quality are commercially available [38]. The QRC method represents an important tool for the realisation of fault-tolerant optical CV quantum computers.

### v3

@@ Abstract
An entirely different calibration method was presented in [13]. It solely utilises squeezed states of light, the Heisenberg uncertainty principle, and relative photon loss measurements. The detection efficiencies of the presumably best photo diodes on the market for 1064 nm and 1550 nm were calibrated to $(99.5\pm 0.5)$% [13] and $(98.5\pm 0.7)$% [14], respectively, using 15-dB and 13.5-dB squeezed vacuum states of light. (In both experiments, a mirror was used to incorporate the 0.5% reflected light from the photodiode during calibration.) The method should be suitable for calibrating photodiodes as primary standards, however, the theoretical description of the quality of the calibration signal is still missing. Furthermore, there is currently no precise measurement method for the photon escape efficiency of the squeezing resonator, which represents a potential source of significant systematic error.
In this work, we present the theoretical description of the calibration signal of quantum radiometric calibration (QRC) as well as the required concept for the measurement of the escape efficiency. We apply the concept in-situ and determine the escape efficiency of our squeezing resonator to (98.583 ± 0.015)%. We use “only” 10-dB squeezed vacuum states, but were nonetheless able to precisely determine the efficiency of the presumably most sensitive photodiodes on the market at 1550 nm (from the same manufacturer as calibrated in [14]) to $\eta_{\rm DE}\!=\!(97.20\pm 0.37)$% and $\eta_{\rm QE}\!=\!(96.9\pm 0.4)$%. Implementing retro-reflection increase the values by 0.46%. The value is unexpectedly low and not sufficient for optical quantum computing.
@@
where the voltages $U$ are transimpedance amplified photo currents measured on continuous-wave light. $U_{\mathrm{tot}}$ is the total measurement voltage and $U_{\mathrm{dark}}$ is the contribution without light. $U_{\mathrm{perf}}$ represents the perfect case, i.e., the unknown voltage that the measurement system would provide if every photon would be converted into exactly one photo electron, combined with zero dark current. In the limit of negligible dark noise, $\eta_{\mathrm{QE}}$ approaches $\eta_{\mathrm{DE}}$. The challenge is to determine accurately $\eta_{\mathrm{DE}}$ and thus $\eta_{\mathrm{QE}}$, i.e., to calibrate a photodetector.
@@
Accordingly, $\eta_{\mathrm{DE}}$ is found when $\eta$ and the other three efficiencies are measured. $\eta_{\mathrm{esc}}$ is the photon escape efficiency from the squeezing resonator, $\eta_{\mathrm{prop}}$ the propagation efficiency from the same resonator to the measurement, and $\eta_{\mathrm{mm}}$ the mode matching efficiency, which corresponds to the square of the interference visibility at the BHD beam splitter. All measurements are photoelectric in nature and take place in situ, i.e., in the same laser-optical QRC setup. This results in low systematic errors and high accuracy. The squeezing resonator’s photon escape efficiency $\eta_{\mathrm{esc}}$ could previously not be measured but just estimated and turned out to be the by far largest error contribution of the method [13]. The concept for measuring $\eta_{\mathrm{esc}}$ is a result of this work and will be discussed later.
We first discuss the measurement principle of $\eta$. Its precision increases with an increasing absolute value of the derivative of the measured uncertainty product to $\eta$. The measured uncertainty product is according to Eq. (2) {align} Δ^2 ^X ⋅Δ^2 ^Y = [ η Δ^2 ^X’ + (1-η) ] ⋅ [ η Δ^2 ^Y’ + (1-η) ]
@@
= 1 + 4(η-η^2) ⟨^n ⟩  , where $(\Delta\!^{2}\!\hat{X}^{\prime}+\Delta\!^{2}\hat{Y}^{\prime})/4-0.5=\langle\hat{n}\rangle$ is the expectation value for the number of quantum correlated photons of the ‘Heisenberg reference’ [20]. For the pure 13.2-dB squeezed state in Fig. 1(b) $\langle\hat{n}\rangle\approx 4.7$.
@@
We introduce the ‘precision scaling’ of QRC $|S_{P}|$, which includes the derivative of Eq. (2) to $\eta$ and the fact that the measurement precisions of $\Delta\!^{2}\!\hat{X}$ and $\Delta\!^{2}\hat{Y}$ are both proportional to the total efficiency $\eta$. {align} —S_P— = η^2  — d (1+ 4(η- η2) ⟨^n⟩)d η — = η^2  —4 - 8 η— ⟨^n ⟩  . Thus, the QRC precision increases with $\langle\hat{n}\rangle$ (whereby a limit is set by the fact that the decoherence due to phase fluctuations must remain negligible, see ‘end matter’).
@@
We applied our QRC to a pair of 1550-nm HQE photodiodes from Laser Components (with removed protection windows), which were optimized for s-polarized continuous-wave laser light at 1550 nm under 20 angle of incidence. Their semiconductor surfaces were free of dust particles. The two photodiodes formed a balanced homodyne detector for measuring $\Delta\!^{2}\!\hat{X}$ and $\Delta\!^{2}\hat{Y}$ of a squeezed state for the determination of $\eta$ according to Eq. (3). These measurements are not performed at direct current, but at a sideband frequency (within the squeezing bandwidth) for which the photodiodes are designated. Here, we chose 5 MHz. We used resonator-enhanced degenerate parametric down-conversion (PDC) in a second-order non-linear crystal for squeezed light generation [22, 23, 24, 13, 31]. A schematic is shown in Fig. 2. The crystals in the SHG and squeezing resonators were periodically poled potassium titanyl phosphate (PPKTP). The crystal geometry was plane-convex and formed a half-monolithic resonator together with a coupling mirror. The crystal was coated to be highly reflective on the convex side and anti-reflective on the other for 1550 nm and 775 nm. The coupling mirror had a design reflectivity of $\approx\!83$ % for 1550 nm and $\approx\!97.5$ % for 775 nm light. The temperature profile of the crystal [32, 33, 34] was stabilized to provide double-resonance as well as good phase matching. The length of the squeezing resonator was stabilized using the Pound-Drever-Hall (PDH) technique [35]. The squeezed vacuum states at 1550 nm (without any carrier light) ‘escaped’ the squeezing resonator through the coupling mirror, were separated from the 775 nm light via a dichroic mirror, and superimposed with the local oscillator on the BHD’s beam splitter. We directly measured a 10-dB squeezed photon shot noise for a 775-nm pump power of slightly less than 11 mW. The anti-squeezing was close to 13 dB. Increasing the pump power increased the squeeze factor just marginally, while the anti-squeezing increased significantly. We used this measurement to estimate the root mean square (rms) phase noise [36] to $\Delta\theta\!<\!(2\!\pm\!1)$ mrad, which was negligible for the QRC result reported further down.
@@
Fig. 3 shows the measurement of $\Delta\!^{2}\!\hat{X}$ (minima) and $\Delta\!^{2}\hat{Y}$ (maxima) for determining the total efficiency $\eta$. Here, we repeatedly swept the BHD read-out phase $\theta$. This approach offers a few advantages over the alternative of locking the read-out phase to each individual quadrature as used before [13, 14]. First of all, it reduces the complexity of the setup by omitting the stabilization of one degree of freedom. Additionally, slow drifts in the local oscillator and pump power can be neglected. Scanning the read-out phase also guarantees that there is no systematic error in the locking point and ensures that each quadrature is optimally measured at some point. The traces in Fig. 3 already include dark noise subtraction and normalisation to the ground state variance. The variance of the ground state was measured with the signal input of the BHD blocked. The variance of the dark noise was measured with all light fields blocked.
@@
The escape efficiency $\eta_{\rm esc}$ of the squeezing resonator could previously only be estimated using an absorption value of the crystal material from the literature [37] and rather uncertain crystal and coupling mirror coating specifications from the manufacturer [13]. Furthermore, potential surface inhomogeneities and small dust particles in the squeezing resonator, which make the value of $\eta_{\rm esc}$ dependent on its alignment, could not be taken into account. In [13], the inaccuracy in the determination of $\eta_{\rm esc}$ clearly dominated the calibration error.
The escape efficiency of the quantum correlated photons from the squeezing resonator is given by
$\eta_{\rm esc}=\frac{1-r^{2}}{1-r^{2}+\ell_{\mathrm{rt}}^{2}}\,$, (7)
where $r^{2}$ is the reflectivity of the ramped coupling mirror in Fig. 4 and $\ell_{\mathrm{rt}}^{2}$ is the cavity’s relative photon loss per roundtrip, with contributions $\ell^{2}_{\mathrm{scat}}$, $\ell^{2}_{\mathrm{abs}}$, $\ell^{2}_{\mathrm{trans}}$, and $\ell^{2}_{\mathrm{refl}}$.
Fig. 4 shows our setup for determining $\eta_{\rm esc}$ (left), a measurement graph of the reflected light power (top right), and the simultaneously measured transmitted power (bottom right). The latter provided a cavity mode matching value of $(98.58\pm 0.23)$% (main peak divided by sum of all peaks), which was required to scale up the loss dip in the top right graph. The fit of Eq. (8) to the up-scaled top right graph required a simultaneous fit of the scanning speed, which we approximated by $\varphi(t)=\varphi_{0}+\varphi_{1}t+\varphi_{2}t^{2}$. Our in-situ measurement yielded $r^{2}=(82.79\pm 0.35)$% and $\ell^{2}_{\mathrm{rt}}=(0.247\pm 0.007)$% and thus $\eta_{\rm esc}=(98.583\pm 0.015)$% according to Eq. (7).
The propagation efficiency $\eta_{\mathrm{prop}}$ and the mode matching efficiency $\eta_{\mathrm{mm}}$ were measured in the established manner, which is described in the ‘end matter’ to this letter. The measurements yielded $\eta_{\mathrm{prop}}\!=(99.49\,\pm\,0.25)$% and $\eta_{\mathrm{mm}}\!=v^{2}=(99.11\,\pm\,0.17)\%$.
@@
The QRC method presented in detail and used here has several unique features compared to established methods. It uses only Heisenberg’s uncertainty principle and the photoelectric effect. Noise powers at the relevant measurement frequency are used for calibration. QRC is thus suitable for determining quantum efficiencies that depend on the measurement frequency. We have shown that absolute calibration with an uncertainty of only 0.37% is possible with squeezed vacuum states in the achievable 10 dB range. Squeeze lasers of this quality are commercially available [38]. The QRC method represents an important tool for the realisation of fault-tolerant optical CV quantum computers.
@@
L.A. performed the final experiments and the data analysis. J.-M.M. performed initial $\eta_{\rm esc}$-measurements. R.S. invented QRC and the $\eta_{\rm esc}$-measurement approach, developed the theoretical QRC description, and prepared the manuscript with the help of L.A.